# Eccentricity evolution of PTA sources from cosmological initial conditions

### v1

@@ 2.3 Computing the orbital parameters
Figure 1: Visual representation of the definition of the fraction $\tilde{f}$, where $r_{\rm{c}1}$ and $r_{\rm{c}2}$ are the centres of mass of the primary and secondary galaxy respectively. The solid blue line represents the enclosed fractional mass profile $f_{1}(\tilde{r}_{1})$ of the primary galaxy, where $\tilde{r}_{1}=r_{\rm{c}1}+r$. For the sake of visual clarity, here the enclosed (fractional) mass profile of the secondary galaxy ($f_{2}(\tilde{r}_{2})$, solid red line) is centred in $r_{\rm{c}2}=r_{\rm{c}1}+r_{\rm{rel}}$ (where $r_{\rm{rel}}$ is the distance between the two centres) and is mirrored with respect to the y-axis (so that the two profiles can intersect); therefore in this representation $\tilde{r}_{2}=r_{\rm{c}2}-r$. The point of intersection $\tilde{r}$ between the two profiles then defines $\tilde{f}$ as $\tilde{f}\equiv f_{1}(\tilde{r})=f_{2}(\tilde{r})$.
@@ 2.4 $N$-body simulations: sub-sample selection and modelling
Figure 2: Density profiles of the primary and secondary galaxy in merger 7, drawn from TNG100-1 (solid black line) and fitted with an Hernquist profile (dashed red line). The black vertical line marks the softening length in the TNG100-1 simulation. First row: DM halo density profiles. Second row: stellar bulge density profiles.
@@ 3.1 Statistics on TNG100-1 mergers
Figure 4: Eccentricity-pericentre parameter space: each black circle represents one of the 127 mergers in our sample. The colour map highlights regions of increasing probability density. The highest density region corresponds to the peak of the distribution, with $e\sim 0.97$ and $r_{\rm{peri}}/r_{\rm{hm,p}}\sim 0.04$.
@@ 3.2 Evolution through binary formation and hardening
Critical times and their respective values of distance and eccentricity are marked in different colours in Fig.5 and 6: (i) the binding time ($t_{\rm{b}}$, red), defined empirically as the time at which the polynomial fit of the eccentricity evolution reaches the minimum. This identifies, in an arbitrary way, the point in the binary evolution when the Keplerian orbital parameters become well defined, and excludes the initial chaotic phase; (ii) the hard-binary time ($t_{\rm{h}}$, blue), defined as the time when the semi-major axis of the binary reaches the hard-binary separation $a_{h}=\frac{q}{(1+q)^{2}}\frac{r_{\rm{m}}}{4}$ (Merritt, 2006), where $q$ is the mass ratio and $r_{\rm{m}}$ is the radius containing a mass in stars equal to twice the mass of the primary; (iii) the time $t_{0,\rm{SAM}}$ where we start the semi-analytical modelling of the evolution (see Sec. 3.3).
@@
Figure 5: Evolution of the distance between the black holes (top panel), Keplerian semi-major axis (middle panel) and eccentricity (bottom panel) in merger 7. The black dashed lines represent the softening parameter $\varepsilon_{0,\rm{BH}}$ (see eq.10) used for the black holes. The coloured dotted lines represent the time $t_{\rm{b}}$ of binary formation (red), the time $t_{\rm{h}}$ when the hard-binary separation is reached (blue), and the time $t_{\rm{0,SAM}}$ when the semi-analytical modelling is started (yellow). The binary forms with very large eccentricity.
@@
Figure 7: Relation between the initial eccentricity of the galactic merger $e_{0}$ and the eccentricity of the binary at binding $e_{\rm{b}}$. We highlight the region with $e_{0}>0.9$ with a red shaded area: above this threshold more circular binaries tend to form, with a significant spread in eccentricity, breaking the expected correlation between $e_{0}$ and $e_{\rm{b}}$
@@ 3.3 Evolution through Gravitational Wave emission and coalescence
In order to highlight the relevance of our results for GWs detection by PTA, we computed the GW frequency of our binaries via $f_{\rm{GW}}=2f_{\rm{orb}}$, where $f_{\rm{orb}}$ is the the orbital frequency [^7] We note that the value of $f_{\rm{GW}}$ thus computed is exact only if the binaries are circular. Moreover, we are not accounting for the redshift, though all of our mergers occur at low redshift, so that its effect is small and can safely be ignored for the sake of our discussion.. In Fig. 11 we plot the orbital parameters as functions of $f_{\rm{GW}}$ and highlight with a red shaded area the frequency band relevant for PTA.
@@
Figure 11: The semi-major axis (top panel) and eccentricity (bottom panel) of the BHBs formed in all of the selected mergers as a function of GW frequency. For visual clarity we plot only the semi-analytical evolution. We highlight the PTA frequency band with a red shaded area.
@@ Acknowledgements
A. Sesana and E. Bortolas acknowledge the financial support provided under the European Union’s H2020 ERC Consolidator Grant “Binary Massive Black Hole Astrophysics” (B Massive, Grant Agreement: 818691). E. Bortolas acknowledges support from the European Union’s Horizon Europe programme under the Marie Skłodowska-Curie grant agreement No 101105915 (TESIFA). The simulations were run on the Eureka2 HPC cluster at the University of Surrey.

### v2

@@ 2.3 Computing the orbital parameters
Figure 1: Visual representation of the definition of the fraction $\tilde{f}$, where $r_{\rm{c}1}$ and $r_{\rm{c}2}$ are the centres of mass of the primary and secondary galaxy respectively. The solid line labelled "Primary" represents the enclosed fractional mass profile $f_{1}(\tilde{r}_{1})$ of the primary galaxy, where $\tilde{r}_{1}=r_{\rm{c}1}+r$. For the sake of visual clarity, here the enclosed (fractional) mass profile of the secondary galaxy ($f_{2}(\tilde{r}_{2})$, solid line labelled "Secondary") is centred in $r_{\rm{c}2}=r_{\rm{c}1}+r_{\rm{rel}}$ (where $r_{\rm{rel}}$ is the distance between the two centres) and is mirrored with respect to the y-axis (so that the two profiles can intersect); therefore in this representation $\tilde{r}_{2}=r_{\rm{c}2}-r$. The point of intersection $\tilde{r}$ between the two profiles then defines $\tilde{f}$ as $\tilde{f}\equiv f_{1}(\tilde{r})=f_{2}(\tilde{r})$.
@@ 2.4 $N$-body simulations: sub-sample selection and modelling
Figure 2: Density profiles of the primary and secondary galaxy in merger 7, drawn from TNG100-1 (solid line) and fitted with an Hernquist profile (dashed line). The vertical line marks the softening length in the TNG100-1 simulation. First row: DM halo density profiles. Second row: stellar bulge density profiles.
@@ 3.1 Statistics on TNG100-1 mergers
Figure 4: Eccentricity-pericentre parameter space: each circle represents one of the 127 mergers in our sample. The colour map highlights regions of increasing probability density. The highest density region corresponds to the peak of the distribution, with $e\sim 0.97$ and $r_{\rm{peri}}/r_{\rm{hm,p}}\sim 0.04$.
@@ 3.2 Evolution through binary formation and hardening
Critical times and their respective values of distance and eccentricity are marked in different colours in Fig.5 and 6 from left to right: (i) the binding time ($t_{\rm{b}}$), defined empirically as the time at which the polynomial fit of the eccentricity evolution reaches the minimum. This identifies, in an arbitrary way, the point in the binary evolution when the Keplerian orbital parameters become well defined, and excludes the initial chaotic phase; (ii) the hard-binary time ($t_{\rm{h}}$), defined as the time when the semi-major axis of the binary reaches the hard-binary separation $a_{h}=\frac{q}{(1+q)^{2}}\frac{r_{\rm{m}}}{4}$ (Merritt, 2006), where $q$ is the mass ratio and $r_{\rm{m}}$ is the radius containing a mass in stars equal to twice the mass of the primary; (iii) the time $t_{0,\rm{SAM}}$ where we start the semi-analytical modelling of the evolution (see Sec. 3.3).
@@
Figure 5: Evolution of the distance between the black holes (top panel), Keplerian semi-major axis (middle panel) and eccentricity (bottom panel) in merger 7. The horizontal dashed lines represent the softening parameter $\varepsilon_{0,\rm{BH}}$ (see eq.10) used for the black holes. The dotted vertical lines represent, from left to right, the time $t_{\rm{b}}$ of binary formation, the time $t_{\rm{h}}$ when the hard-binary separation is reached, and the time $t_{\rm{0,SAM}}$ when the semi-analytical modelling is started. The binary forms with very large eccentricity.
@@
Figure 7: Relation between the initial eccentricity of the galactic merger $e_{0}$ and the eccentricity of the binary at binding $e_{\rm{b}}$. We highlight the region with $e_{0}>0.9$ with a shaded area: above this threshold more circular binaries tend to form, with a significant spread in eccentricity, breaking the expected correlation between $e_{0}$ and $e_{\rm{b}}$
@@ 3.3 Evolution through Gravitational Wave emission and coalescence
In order to highlight the relevance of our results for GWs detection by PTA, we computed the GW frequency of our binaries via $f_{\rm{GW}}=2f_{\rm{orb}}$, where $f_{\rm{orb}}$ is the the orbital frequency [^7] We note that the value of $f_{\rm{GW}}$ thus computed is exact only if the binaries are circular. Moreover, we are not accounting for the redshift, though all of our mergers occur at low redshift, so that its effect is small and can safely be ignored for the sake of our discussion.. In Fig. 11 we plot the orbital parameters as functions of $f_{\rm{GW}}$ and highlight with a shaded area the frequency band relevant for PTA.
@@
Figure 11: The semi-major axis (top panel) and eccentricity (bottom panel) of the BHBs formed in all of the selected mergers as a function of GW frequency. For visual clarity we plot only the semi-analytical evolution. We highlight the PTA frequency band with a shaded area.
@@ Acknowledgements
AS and EB acknowledge the financial support provided under the European Union’s H2020 European Research Council (ERC) Consolidator Grant “Binary Massive Black Hole Astrophysics” (B Massive, Grant Agreement: 818691). AG would like to thank the Science and Technology Facilities Council (STFC) for support from grant ST/Y002385/1. EB acknowledges support from the European Union’s Horizon Europe programme under the Marie Skłodowska-Curie grant agreement No 101105915 (TESIFA). The simulations were run on the Eureka2 HPC cluster at the University of Surrey.